OpenStudy (mef4.0):

How do you find the derivative of: f(s) = cot x/(1+csc x)

OpenStudy (slaaibak):

I think the quotient rule would be better

OpenStudy (slaaibak):

http://en.wikipedia.org/wiki/Quotient_rule

OpenStudy (mertsj):

bottom times the derivative of the top - top times derivative of bottom all over bottom squared

OpenStudy (anonymous):

this is an annoying problem. change it first to \[\frac{\cot(x)}{1+\csc(x)}=\frac{\cos(x)}{\sin(x)+1}\] and take the derivative of that with the quotient rule

OpenStudy (anonymous):

that one is easy \[\frac{-(\sin(x)+1)\times \sin(x)-(\cos^2(x))}{(\sin(x)+1)^2}\] \[\frac{-\sin^2(x)-\sin(x)-\cos^2(x)}{(\sin(x)+1)^2}\] \[\frac{-\sin(x)-1}{(\sin(x)+1)^2}\] \[-\frac{1}{\sin(x)+1}\]

OpenStudy (mef4.0):

thank you